# BERP: A Blind Estimator of Room Parameters for Single-Channel Noisy Speech Signals

### v6

@@ Abstract
Room acoustical parameters (RAPs), room geometrical parameters (RGPs) and instantaneous occupancy level are essential metrics for parameterizing the room acoustical characteristics (RACs) of a sound field around a listener’s local environment, offering comprehensive indications for various applications. Current blind estimation methods either fail to cover a broad range of real-world acoustic environments in the context of real background noise or estimate only a few RAPs and RGPs from noisy single-channel speech signals. In addition, they are limited in their ability to estimate the instantaneous occupancy level. In this paper, we propose a new universal blind estimation framework called the blind estimator of room parameters (BERP) to estimate RAPs, RGPs and occupancy level via a unified methodology. It consists of two modules: a unified room feature encoder that combines attention mechanisms with convolutional layers to learn common features across room parameters, and multiple separate parametric predictors for continuous estimation of each parameter in parallel. The combination of attention and convolutions enables the model to capture acoustic features locally and globally from speech, yielding more robust and multitask generalizable common features. Separate predictors allow the model to independently optimize for each room parameter to reduce task learning conflict and improve per-task performance. This estimation framework enables universal and efficient estimation of room parameters while maintaining satisfactory performance. To evaluate the effectiveness of the proposed framework, we compile a task-specific dataset from several publicly available datasets, including synthetic and real reverberant recordings. The results reveal that BERP achieves state-of-the-art (SOTA) performance and excellent adaptability to real-world scenarios. The code, datasets, and weights are available on GitHub 22footnotemark: 2.
@@
Subjective listening tests can be conducted to assess local RACs in a listener’s surroundings. However, conducting a subjective test is expensive and time-consuming, and more importantly, subjective results are significantly influenced by individual differences such as age, hearing ability, and cultural background. We usually instead use room acoustical parameters (RAPs) as the objective indices to assess the local RACs, such as the reverberation time and the speech intelligibility index. Room geometrical parameters (RGPs), such as room volume, sound source distance and direction of arrival (DOA), are also important for evaluating local RACs, as they have critical applications in spatial audio rendering, intelligibility assessments in a room, separation of sound sources, audio navigation systems and speech enhancement [1, 2, 3, 4, 5, 6, 7]. Consequently, RAPs and RGPs can be used together to model local RACs to offer clear and comprehensive indications for various applications, such as room acoustical assessment [8, 9, 10, 5, 11], speech enhancement [12, 13, 14], hearing aids [15, 16, 17, 18, 4, 14], sound source separation [1, 2], spatial audio rendering [19], context-aware rendering in extended reality (XR) and augmented reality (AR) [20, 3, 21], public address systems [22, 23], and robotics [24]. Another important metric is instantaneous occupancy level, which quantifies how many people are present near the listener at a given time; this is an essential consideration in demand-driven HVAC (heating, ventilation and air conditioning) control systems for smart buildings [25, 26].
@@
Several RAPs have been investigated and standardized [27, 9, 8, 10], including the speech transmission index (STI), the percentage articulation loss of consonants () reverberation time (), early decay time (EDT), clarity ( / ), definition () and center time (). RAPs can be directly derived from the measured RIR. However, measuring RIR requires the exclusion of people in an enclosure, which is impractical for public spaces, as the RIR measurement requires high-energy sounds[28, 29]. Furthermore, it is limited in capturing the dynamics of local RACs, which vary according to the locations, arrangements, and quantities of objects and occupants that are present. The RAPs measured using specific standards may differ from noncompliant measures employed within the same enclosure. Meanwhile, RGPs may be derived from the measured RIR, but they encounter the same aforementioned issues. Therefore, some blind estimation methods have been proposed to estimate RAPs and RGPs, particularly in public spaces where people cannot be excluded.
@@
In general, blind estimation is a challenging task because it is an ill-posed inverse problem without prior knowledge about systems and inputs. In this case, conventional blind estimation methods use mathematical derivations to create an explicit map between observed signals and systems[30, 31, 28], whereas current blind estimation methods typically employ deep neural networks (DNNs) to map observed signals to room parameters due to their superior performance over conventional analytical methods[32], such as convolutional neural networks (CNNs) [33, 34, 35, 36, 37], convolutional recurrent neural networks (CRNNs) [38, 39, 40, 41], and hybrid approaches[29].
@@
Despite these advances, two key limitations remain. First, most methods estimate only a small subset of parameters (e.g., one, two, or four parameters), which is insufficient for fully assessing local RACs. None of them can simultaneously predict the full set of RAPs and RGPs. Although hybrid approaches can theoretically handle multiple parameters at once, their accuracy is limited by how well the RIR model matches real-world conditions[29], and they fail to address RGPs. In this direction, we propose a solution to address this issue, i.e., blind estimation of a full set of RAPs and RGPs. We formulate this task as a multitask learning (MTL) problem, treating each room parameter as a separate task. RAPs and RGPs have mutual relatedness, as they share common reverberation and geometric information within the RIR. Cerdá used factor analysis to analyze correlations among RAPs[42] and studies on Sabine’s formula and the critical distance of intelligibility () revealed a correlation between room volume and reverberation time[27, 4]. [4] also explored the relationships among , room volume, and sound source distance. Figure 1 illustrates the mutual relatedness among the room parameters. On the one hand, exploiting mutual relatedness among room parameters may improve per-task performance, particularly for harder tasks or tasks with limited training examples, since task relatedness promotes knowledge transfer to allow more effective learning [43, 44]. On the other hand, adding more room parameters as related tasks provides auxiliary task learning benefits to further improve the performance of one or more primary tasks, especially when those tasks have no strong relatedness[44, 45]. At the same time, MTL often suffers from task learning conflicts[45]. One possible solution is to focus on model design to reduce task training conflicts and improve multitask performance.
@@
Second, a key challenge lies in blind estimation of instantaneous occupancy level. While RAP and RGP estimation have received considerable attention, the estimation of instantaneous occupancy level has been largely overlooked; this may be because RAP and RGP estimation typically assume a single-source speech signal with background noise, whereas the occupancy level estimation assumes a complex, overlapping speech signal, which seems to be significantly more challenging. To simplify the task, some existing methods treat occupancy as a stationary status to estimate a single value[25, 46]. However, occupancy in the real world changes dynamically over time, making it more appropriate to predict a time sequence. Moreover, they may not provide a meaningful reflection of performance, as F1 score regards any prediction that is not exactly correct as entirely wrong, which may not capture how close a predicted count is to the true count. Instead, directly measuring how closely the predicted count aligns with the ground truth by using distance metrics (e.g., mean absolute error (MAE)) is much better. To our knowledge, predicting instantaneous occupancy levels over time, particularly considering room acoustical settings, has received little attention. If possible, we expect this task to be unified with RAP and RGP estimation in a universal framework. Owing to the scarcity of publicly available data for instantaneous occupancy-level estimation, we introduce a data preparation approach specifically designed to address this gap.
@@
Motivated by these findings, we propose a new framework (BERP) that can jointly estimate a full set of room parameters and the occupancy level while preserving satisfactory per-task performance, as shown in Fig. 2. More specifically, we design a modular architecture that consists of a shared encoder and multiple task-specific predictors for the joint estimation of RAPs and RGPs. Capturing acoustic features locally and globally together from noisy reverberant speech with complicated waveform may enable the encoder to learn more robust and multitask-generalizable common features since acoustical information spreads the observed speech signal[28]. Therefore, the shared encoder combines attention and CNNs to grasp features in a time-frequency representation both globally and locally[47, 48]. The task-specific predictor, acting as a regressor, predicts a continuous estimation of a room parameter from learned common features.
Regarding the blind estimation of the occupancy level, inspired by automatic speech recognition that classifies tokens in each time frame, we formulate this task as a sequence classification problem. Because this problem involves time series data and given that the aforementioned encoder excels in extracting features for a sequence, we directly apply the same encoder architecture. Reusing the encoder, we estimate RAPs, RGPs, and the occupancy level within the same framework.
@@
Currently, except for the occupancy level, BERP assumes a single-source speech with background noise as the observed signal. This assumption, while effective for many scenarios, may not fully capture the complexity of real-world acoustic environments. More discussion of current limitations and future work is included in Section VII.
@@
A new universal blind estimation framework is proposed to blindly estimate room parameters, which achieves state-of-the-art (SOTA) performance
@@
The remainder of the manuscript is organized as follows. Section II provides literature reviews of state-of-the-art works. Section III briefly reviews the room parameters, including the RAPs, RGPs, and instantaneous occupancy level. Section IV specifies the data generation pipeline for the proposed method. The proposed approach is introduced in Section V. The experiments are described in Section VI. We discuss and conclude in Section VII.
@@ II Related works
Further advances include combining CNNs with MLPs to jointly estimate , volume, and room surface using single- or two-channel noisy speech [35], as well as employing gating mechanisms within CNNs for estimation [34]. Duangpummet et al.[29] proposed a hybrid model that integrates RIR modeling with CNNs, allowing concurrent estimation of STI, , , , and . They first approximate the RIR with an RIR model to estimate the parameters of this model and then derive room parameters from the synthesized RIR. Attention have also gained traction: Wang et al.[49, 50] explored pure attention mechanisms for estimating and volume. Meanwhile, new training strategies are emerging beyond supervised end-to-end approaches. Philipp et al.[51] introduced a task-agnostic embedding method to replace direct supervised learning.
@@
Most current methods can estimate only a small subset of room parameters, which is insufficient for comprehensively describing local RACs. We argue that this limitation may result from their architectural design. They use a single architecture in which all tasks share the same neural network parameters and produce multiple room parameters from one output layer. This setup may lead to serious task-training conflicts and undermine per-task performance, since each task is forced to compete within the same optimization space. One task’s gradient update may severely disrupt another, causing gradient conflicts and making it difficult for the model to find an effective optimization path that satisfies all tasks from the start[52]. As a result, when estimating many parameters (such as ten), more serious task conflicts may significantly undermine their potential to obtain satisfactory per-task performance.
@@
Inspired by these explorations, we develop a new model. In essence, our method differs algorithmically in how it manages shared and task-specific optimizations to improve multitask performance and reduce task conflicts[53, 45]. For this purpose, we decompose the model into two parts: a shared encoder that captures common features and multiple task-specific predictors for given room parameters. This architecture reduces interference by preventing multiple tasks from vying for the same output layer, which enables each predictor specializing in its task while encoder focus on learning common features more beneficial across all tasks. More specifically, we combine attention and convolutions to design the shared encoder. The task-specific predictor consists of convolutions that are well suited for the regression task. There are three key motivations behind this design. First, attention enables our encoder to capture dependencies across any two positions, incorporating both long-range and global contextual information; this is crucial in multitask settings, where different tasks may require attention to various segments of the input. This encoder can extract high-level features to feed into multiple predictors for different tasks, reducing parameter redundancy and fostering knowledge sharing between tasks[54]. In contrast, traditional CNNs or RNNs used in baselines struggle to capture such long-range dependencies and are less adaptable to flexible multitask outputs. Next, using a shared encoder provides an inductive bias that learns valuable features across multiple tasks, promoting both knowledge transfer and a natural regularization effect to reduce overfitting. Past studies have shown that such inductive biases lead to more robust task-generalizable models [55, 45, 44, 54]. Third, this architecture leverages the strength of convolutions in extracting local features while simultaneously using attention for the global context [47]. This design is advantageous in this case, where both local nuances and the global structure are important for learning acoustic features, since acoustic information smears all time-frequency components of reverberant inputs[28].
@@
Another aspect is training data. Many current methods rely on synthetic RIRs via RIR simulation algorithms because they can be easily generated on a large scale [36, 34, 33, 29]. Common simulation algorithms include FDTD-based approaches such as ARD [56] used in [33]; image source-based methods such as pyroomacoustic[57] used in[33] and an RIR Generator[58] used in [34]; and a stochastic approach [29]. However, synthetic RIRs generated from these simulation algorithms have limitations. RIRs from image source methods often simplify complex geometries and materials with sophisticated absorption and scattering coefficients and wave phenomena (e.g., diffraction) due to intricate wall material, decorations, and furniture in real rooms, to reduce computational overhead [59, 60, 33], although source image principle is suitable for modeling RIR. FDTD-based ARD also lacks certain real-world irregularities [36], risking overfitting when large synthetic datasets are used [61]. Similarly, the stochastic method [29] struggles to replicate the sparse onset properties of real RIRs [62]. Building on these insights, we generate synthetic data exclusively from real RIRs for blind estimation of RAPs and RGPs. To enhance adaptability to real-world conditions, we supplement our large synthetic dataset with a smaller set of real recordings. While synthetic data offer a large and easily controlled training environment to help the model quickly learn core patterns and stabilize training, real records capture nuanced characteristics of real settings. This data preparation method is given that relying solely on real-world annotated data is challenging, as such datasets are often scarce and costly to collect at a large scale.
@@
Besides, considering that the observed signal is restricted to a single-channel omnidirectional format that cannot sufficiently distinguish DOA without known room geometry (blind in our case)[63], we exclude DOA from our target RGPs.
@@
Blind estimation of occupancy level. In contrast to RAPs and RGPs, blind occupancy level estimation has received far less attention. The few existing studies treat occupancy as a static classification problem to predict a single value rather than a time sequence[25, 46]. However, in real-world settings, occupancy fluctuates over time, making instantaneous occupancy-level estimation more appropriate. Moreover, these methods ignore room acoustics, which significantly influences speech transmission. Another issue arises from their evaluation method. They use F1 score rather than quantifying the distance between predicted and actual occupancy levels (e.g., MAE), which may not provide a meaningful reflection of performance. For example, predicting 9 occupants instead of 10 is much more appropriate than predicting 2, yet both are simply “incorrect” under a strict accuracy measure.
@@
Given these issues, we therefore redefine the task as instantaneous occupancy level estimation, formulating it as a classification problem at each time frame while considering room acoustics. To accomplish this, we design a dedicated model that includes an encoder and a classifier head. In addition, owing to the scarcity of publicly available data for instantaneous occupancy level estimation, we introduce a data preparation approach specifically designed to address this gap. The corresponding data generation pipeline is detailed in Section IV. Besides, instead of F1 score, we use the macro-averaged MAE and MAE to evaluate the model’s performance.

### v7

@@ Abstract
Room acoustical parameters (RAPs), room geometrical parameters (RGPs) and instantaneous occupancy levels are essential metrics for parameterizing the room acoustical characteristics (RACs) of a sound field around a listener’s local environment, offering comprehensive indications for various applications. Current blind estimation methods either fail to cover a broad range of real-world acoustic environments in the context of real background noise or estimate only a few RAPs and RGPs from noisy single-channel speech signals. In addition, they are limited in their ability to estimate the instantaneous occupancy level. In this paper, we propose BERP, a new universal approach to blindly estimate RAPs, RGPs, and occupancy levels. It consists of two modules: one for RAPs and RGPs and another for occupancy levels. For the former task, we use a shared room feature encoder that combines attention mechanisms with convolutional layers to learn common features across room parameters, and multiple separate parametric predictors for continuous estimation of each parameter in parallel. The combination of attention and convolutions enables the model to capture acoustic features both locally and globally from speech, yielding more robust and multitask generalizable common features. Separate predictors allow the model to independently optimize for each room parameter to reduce task learning conflict and improve per-task performance. This architecture enables universal and efficient estimation of room parameters while maintaining satisfactory performance. For occupancy level estimation, we reuse the identical encoder with a classification head, exploiting the encoder’s strong sequence‑feature extraction. To evaluate the effectiveness of the proposed approach, we compile a task-specific dataset from several publicly available datasets, including synthetic and real reverberant recordings. The results reveal that BERP achieves state-of-the-art (SOTA) performance and excellent adaptability to real-world scenarios. The code, datasets, and weights are available on GitHub 22footnotemark: 2.
@@
Subjective listening tests can be conducted to assess local RACs in a listener’s surroundings. However, conducting a subjective test is expensive and time-consuming, and more importantly, subjective results are significantly influenced by individual differences such as age, hearing ability, and cultural background. Instead, room acoustical parameters (RAPs), such as the reverberation time and the speech intelligibility index, are usually used as the objective indices to assess the local RACs. Room geometrical parameters (RGPs), such as the room volume, sound source distance and direction of arrival (DOA), are also important for evaluating local RACs, as they have critical applications in spatial audio rendering, intelligibility assessments in a room, separation of sound sources, audio navigation systems and speech enhancement [1, 2, 3, 4, 5, 6, 7]. Consequently, RAPs and RGPs can be used together to model local RACs to offer clear and comprehensive indications for various applications, such as room acoustical assessment [8, 9, 10, 5, 11], speech enhancement [12, 13, 14], hearing aids [15, 16, 17, 18, 4, 14], sound source separation [1, 2], spatial audio rendering [19], context-aware rendering in extended reality (XR) and augmented reality (AR) [20, 3, 21], public address systems [22, 23], and robotics [24]. Another important metric is the instantaneous occupancy level, which quantifies how many people are present near the listener at a given time; this is an essential consideration in demand-driven HVAC (heating, ventilation, and air conditioning) control systems for smart buildings [25, 26].
@@
Several RAPs have been investigated and standardized [27, 9, 8, 10], including the speech transmission index (STI), the percentage articulation loss of consonants (), reverberation time (), early decay time (EDT), clarity ( / ), definition () and center time (). RAPs can be directly derived from the measured RIR. However, measuring RIR requires the exclusion of people in an enclosure, which is impractical for public spaces, as it involves the use of high-energy sounds[28, 29]. Furthermore, it is limited in capturing the dynamics of local RACs, which vary depending on the locations, arrangements, and quantities of objects and occupants present. The RAPs measured using specific standards may differ from noncompliant measures employed within the same enclosure. Moreover, RGPs may be derived from the measured RIR, but they encounter the same aforementioned issues. Therefore, some blind estimation methods have been proposed to estimate RAPs and RGPs, particularly in public spaces where people cannot be excluded.
@@
In general, blind estimation is a challenging task because it is an ill-posed inverse problem without prior knowledge about systems and inputs. In this case, conventional blind estimation methods use mathematical derivations to create an explicit map between observed signals and systems[30, 31, 28], whereas current blind estimation methods typically employ deep neural networks (DNNs) to map observed signals to room parameters owing to their superior performance over conventional analytical methods[32], such as convolutional neural networks (CNNs) [33, 34, 35, 36, 37, 38], convolutional recurrent neural networks (CRNNs) [33, 39, 40, 41, 42], and hybrid approaches[29].
@@
Despite these advances, two key limitations remain. First, most methods estimate only a small subset of parameters (e.g., one, two, or four parameters), which is insufficient for fully assessing local RACs. None of them can simultaneously predict the full set of RAPs and RGPs. Although hybrid approaches can theoretically handle multiple parameters at once, their accuracy is limited by how well the RIR model matches real-world conditions[29], and they fail to address RGPs. In this direction, we propose a solution to address this issue, i.e., blind estimation of a full set of RAPs and RGPs. We formulate this task as a multitask learning problem, treating each room parameter as a separate task. RAPs and RGPs have mutual relatedness, as they share common reverberation and geometric information within the RIR. Cerdá analyzed the correlations among RAPs[43] and studies on Sabine’s formula and the critical distance of intelligibility () revealed a correlation between room volume and reverberation time[27, 4]. [4] also explored the relationships among , room volume, and sound source distance. Figure 1 illustrates the mutual relatedness among the room parameters as shown in [43, 27], and [4]. On the one hand, exploiting mutual relatedness among room parameters may improve per-task performance, particularly for harder tasks or tasks with limited training examples, since task relatedness promotes knowledge transfer to allow more effective learning [44, 45]. On the other hand, adding more room parameters as related tasks provides auxiliary task learning benefits to further improve the performance of one or more primary tasks, especially when those tasks have no strong relatedness[45, 46]. Moreover, it often suffers from task learning conflicts[46]. One possible solution is to focus on model design to reduce task training conflicts and improve multitask performance.
@@
Second, a key challenge lies in the blind estimation of the instantaneous occupancy level. While RAP and RGP estimations have received considerable attention, estimation of the instantaneous occupancy level has been largely overlooked; this may be because RAP and RGP estimations typically assume a single-source speech signal with background noise, whereas the occupancy level estimation assumes a complex, overlapping speech signal, which seems to be significantly more challenging. To simplify the task, some existing methods treat occupancy as a stationary status to estimate a single value[25, 47]. However, occupancy in the real world changes dynamically over time, making it more appropriate to predict a time sequence. Moreover, these methods may not provide a meaningful reflection of performance, as the F1 score considers any prediction that is not exactly correct as entirely wrong, which may not capture how close a predicted count is to the true count. Instead, directly measuring how closely the predicted count aligns with the ground truth by using distance metrics (e.g., mean absolute error (MAE)) is much more precise. To our knowledge, predicting instantaneous occupancy levels over time, particularly considering room acoustic settings, has received little attention. If possible, we expect this task to be unified with RAP and RGP estimations via a universal methodology. Owing to the scarcity of publicly available data for instantaneous occupancy-level estimation, we introduce a data preparation approach specifically designed to address this gap.
@@
Motivated by these findings, we propose a new approach (BERP) that tackles two distinct tasks: (1) joint estimation of RAPs and RGPs (namely unified module) and (2) estimation of occupancy levels (namely occupancy module), as shown in Fig. 2. More specifically, for the joint estimation of RAPs and RGPs, we design a modular architecture that consists of a shared encoder and multiple task-specific predictors. Capturing acoustic features locally and globally together from noisy reverberant speech with complicated waveforms may enable the encoder to learn more robust and multitask-generalizable common features since acoustical information spreads the observed speech signal[28]. Therefore, the shared encoder combines attention and CNNs to grasp features in a time-frequency representation both globally and locally[48, 49]. The task-specific predictor, which acts as a regressor, predicts a continuous estimation of a room parameter from learned common features. Regarding the blind estimation of the occupancy level, inspired by automatic speech recognition that classifies tokens in each time frame, we formulate this task as a sequence classification problem. Because this problem involves time series data and given that the aforementioned encoder excels in extracting features for a sequence, we directly apply the same encoder architecture. Reusing the encoder, we estimate RAPs, RGPs, and the occupancy level within the same methodology.
@@
Currently, except for the occupancy level, BERP assumes a single-source speech signal with background noise as the observed signal. This assumption, while effective for many scenarios, may not fully capture the complexity of real-world acoustic environments. More discussion of current limitations and future work is included in Section VII.
@@
A new universal blind estimation approach is proposed to blindly estimate room parameters and occupancy levels, which achieves state-of-the-art (SOTA) performance.
@@
The remainder of the manuscript is organized as follows. Section II provides a literature reviews of state-of-the-art works. Section III briefly reviews the room parameters, including the RAPs, RGPs, and instantaneous occupancy level. Section IV specifies the data generation pipeline for the proposed method. The proposed approach is introduced in Section V. The experiments are described in Section VI. We discuss the findings and conclude in Section VII.
@@ II Related works
Further advances include combining CNNs with MLPs to jointly estimate , volume, and room surface using single- or two-channel noisy speech [36] and employing gating mechanisms within CNNs for estimation [35]. Duangpummet et al.[29] proposed a hybrid model that integrates RIR modeling with CNNs, allowing concurrent estimation of STI, , , , and . Attention has also gained traction: Wang et al.[50, 51] explored pure attention mechanisms for estimating and volume. Meng et al.[33] estimated and for ambisonics recordings using the CNN and the CRNN architectures. Moreover, new training strategies are emerging beyond supervised end-to-end approaches. Philipp et al.[52] introduced a task-agnostic embedding method to replace direct supervised learning.
@@
Most current methods can estimate only a small subset of room parameters, which is insufficient for comprehensively describing local RACs. We argue that this limitation may result from their architectural design. They use a single architecture in which all tasks share the same neural network parameters and produce multiple room parameters from one output layer. This setup may lead to serious task-training conflicts and undermine per-task performance, since each task is forced to compete within the same optimization space. One task’s gradient update may severely disrupt another, causing gradient conflicts and making it difficult for the model to find an effective optimization path that satisfies all tasks from the start[53]. As a result, when estimating many parameters (such as ten), more serious task conflicts may significantly undermine their potential to achieve satisfactory per-task performance.
@@
Inspired by these explorations, we develop a new model. In essence, our method differs algorithmically in how it manages shared and task-specific optimizations to improve multitask performance and reduce task conflicts[54, 46]. For this purpose, we decompose the model into two parts: a shared encoder that captures common features and multiple task-specific predictors for given room parameters. This architecture reduces interference by preventing multiple tasks from vying for the same output layer, which enables each predictor to specialize in its task, while the encoder focuses on learning common features that are more beneficial across all tasks. More specifically, we combine attention and convolutions to design the shared encoder. The task-specific predictor consists of convolutions that are well suited for the regression task. There are three key motivations behind this design. First, attention enables our encoder to capture dependencies across any two positions, incorporating both long-range and global contextual information; this is crucial in multitask settings, where different tasks may require attention to various segments of the input. This encoder can extract high-level features to feed into multiple predictors for different tasks, reducing parameter redundancy and fostering knowledge sharing between tasks[55]. In contrast, traditional CNNs or RNNs used in baselines struggle to capture such long-range dependencies and are less adaptable to flexible multitask outputs. Next, using a shared encoder provides an inductive bias that learns valuable features across multiple tasks, promoting both knowledge transfer and a natural regularization effect to reduce overfitting. Past studies have shown that such inductive biases lead to more robust task-generalizable models [56, 46, 45, 55]. Third, this architecture leverages the strength of convolutions in extracting local features while simultaneously using attention for the global context [48]. This design is advantageous in this case, where both local nuances and the global structure are important for learning acoustic features since acoustic information smears all time-frequency components of reverberant inputs[28].
@@
Another aspect is the training data. Many current methods rely on synthetic RIRs via RIR simulation algorithms because they can be easily generated on a large scale [37, 35, 34, 29]. Common simulation algorithms include FDTD-based approaches such as ARD [57] used in [34]; image source-based methods such as pyroomacoustic[58] used in[34] and an RIR Generator[59] used in [35]; and a stochastic approach [29]. However, synthetic RIRs generated from these simulation algorithms have limitations. RIRs from image source methods often simplify complex geometries and materials with sophisticated absorption and scattering coefficients, as well as wave phenomena (e.g., diffraction), due to intricate wall materials, decorations, and furniture in real rooms, to reduce computational overhead[60, 61, 34], although source image principle is suitable for modeling RIR. FDTD-based ARD also lacks certain real-world irregularities[37], risking overfitting when large synthetic datasets are used [62]. Similarly, the stochastic method [29] struggles to replicate the sparse onset properties of real RIRs [63]. Building on these insights, we generate synthetic data exclusively from real RIRs for blind estimation of RAPs and RGPs. To enhance adaptability to real-world conditions, we supplement our large synthetic dataset with a smaller set of real recordings. While synthetic data offer a large and easily controlled training environment to help the model quickly learn core patterns and stabilize training, real records capture nuanced characteristics of real settings. This data preparation method is given that relying solely on real-world annotated data is challenging, as such datasets are often scarce and costly to collect at a large scale.
@@
In addition, considering that the observed signal is restricted to a single-channel omnidirectional format that cannot sufficiently distinguish DOA without known room geometry (blind in our case)[64], we exclude DOA from our target RGPs.
@@
Blind estimation of occupancy level. In contrast to RAPs and RGPs, blind occupancy level estimation has received far less attention. The few existing studies treat occupancy as a static classification problem to predict a single value rather than a time sequence[25, 47]. However, in real-world settings, occupancy fluctuates over time, making instantaneous occupancy-level estimation more appropriate. Moreover, these methods ignore room acoustics, which significantly influences speech transmission. Another issue arises from their evaluation method. These methods use the F1 score rather than quantifying the distance between the predicted and actual occupancy levels (e.g., MAE), which may not provide a meaningful reflection of performance. For example, predicting 9 occupants instead of 10 is much more appropriate than predicting 2, yet both are simply “incorrect” under a strict accuracy measure.
@@
Given these issues, we therefore redefine the task as instantaneous occupancy level estimation, formulating it as a classification problem at each time frame while considering room acoustics. To accomplish this, we design a dedicated model that includes an encoder and a classifier head. In addition, owing to the scarcity of publicly available data for instantaneous occupancy level estimation, we introduce a data preparation approach specifically designed to address this gap. The corresponding data generation pipeline is detailed in Section IV. Besides, instead of the F1 score, we use the macro-averaged MAE and MAE to evaluate the model’s performance.